OpenStudy (anonymous):

what is the measure of one interior angle in a regular hexagon? 360 36 108 120

OpenStudy (anonymous):

The measure of any one angle in a regular polygon is equal to 180(n-2)/n, where n is the amount of sides of the figure. So a hexagon has 6 sides. Thus we plug in 6 for n \[\frac{ 180(6-2) }{ 6 } = \frac{ 180(4) }{ 6 } = 30(4) = 120\]